OpenStudy (anonymous):

Expand the squared binomial. (x 4)2

OpenStudy (anonymous):

what is multiplied by 4?. Use binomial theorem ( ) ( ) ( ) ( ) N where ! !( )!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

if it is square then formula is (a+b)^2=a^2+2*a*b+b^2

OpenStudy (anonymous):

x2 16???

OpenStudy (anonymous):

+

OpenStudy (anonymous):

\[(a+b)^{n}=a^{n}+\left(\begin{matrix}n \\ 1\end{matrix}\right)a ^{n-1}b+...till +b\]

OpenStudy (anonymous):

b is in "n" degree

OpenStudy (anonymous):

I came up with ---- x2 + 8x 16

OpenStudy (anonymous):

Am I doing something wrong?

OpenStudy (anonymous):

is "x" a variable or sign of multiplication?

OpenStudy (anonymous):

is 2 is square?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Multiplication.

OpenStudy (anonymous):

then what the number is placed before "x"?

OpenStudy (anonymous):

2 squared.

OpenStudy (anonymous):

\[(2^{2}*4)^{2}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

No, (x-4)2

OpenStudy (anonymous):

then, \[x^{2}-2*x*4+4^{2}\] =>> \[x^{2}-8x+16\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Yep, Thanks.